1. Mennen Corporation has 30,000 shares of $10 par value common stock outstanding when it announces a 2-for-1 stock split. Before the split, the stock had a market price of $120 per share. After the split, how many shares of stock will be outstanding? What will be the approximate market price per share?


1
Expert's answer
2021-03-21T20:05:24-0400
"30000+\frac{30000}{1}\times 2=90000"

"\frac {120\times 30000}{90000}=40"